I need step by step solution to the following this question asap .I have limited time so please do it quickly with detailed explanation

thanks in advance/Ha

( poms) Question 3 Consider the market for the homogenous good space dust with the following inverse demand function: p(y)
Answer #1

Answer: 3

Answey Here we have used the concepts oelated the Monopoly and digopoly Market structures with the increase in the number ofDY © Tatal Revenue ob Samuotse beeline is TR =P OYJ = 822 - 289}= 22Y - 272 i profit , It is given by T - Tatal Revenue - TotLys 22 -484 x = 4.5 Y = 4.5 litres will be the quantity of beer sold: The market - P(y) =18 units will be the market priceTg = P(y)Y, - YY, >> Ti =(22-27, -272) Y1 - 470 dAcatling ito Frooc. - 22-471-242-420 3) 18 - 471 -2y2 = 0 => Y, = 18 - 272.Ra. () is the best response function of frisyne (inbeyms y yu) LRL%and Ky () aye The best blesponse functions of film i and f4 Pulling the value of Ye fejom ☺ in D, we gets- Y, - 18-24%) = 36- 18 +241 - y = 10th = 9+1 - - 341 - 9 al = 3. => Y = 3 lid